OpenStudy (anonymous):

how to solve this equation: √3-x=x-1

OpenStudy (callisto):

is this your question? \[\sqrt{3-x}=x-1\]

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[\sqrt{3-x}=x-1\] 1. square both sides \[\sqrt{3-x}^2=(x-1)^2\]\[3-x=x^2-2x+1\]\[x^2-x-2=0\] Can you solve it now?

OpenStudy (anonymous):

that's as far as i got i didnt know what to do there?

OpenStudy (callisto):

Solving quadratic equation... either factorise it or applying the quadratic formula..

OpenStudy (anonymous):

oh okay i got x=2 and x=-1

OpenStudy (callisto):

Yes !~

OpenStudy (callisto):

it should be ''or'' not ''and''

OpenStudy (anonymous):

it's x=2

OpenStudy (callisto):

Oh... I forgot a step. You need to put the answer back into the equation and check it. If the right hand side has a negative value, you need to reject it!

OpenStudy (callisto):

So, x=-1 would not be the answer :|